Calculate the pressure exerted on a diver at 60 m below the free surface of the sea.

The barometric pressure is 101 kPa and the specific gravity of seawater is 1.03.


1
Expert's answer
2022-06-01T12:49:30-0400
"p=p_0+g\rho h\\
=101*10^3+1.03*10^3*9.8*60=7.07*10^5\:\rm Pa"